OpenStudy (snowystorm777):

How many solutions does the following equation have? |3x + 12| = 18

OpenStudy (mathmale):

I think you should actually try to find the solution or solutions before trying to answer this question. You could simplify the problem considerably by dividing every term by 3.

OpenStudy (snowystorm777):

so x+4=6?

OpenStudy (mathmale):

this is an absolute value problem, so it needs to be treated a bit differently. You now have |x+4|=6. How many solutions does that have? What does the presence of the absolute value operator, | | , tell you to do?

OpenStudy (anonymous):

\[\left| x+a \right|=b,x+a=\pm b\]

OpenStudy (snowystorm777):

make it positive?

OpenStudy (snowystorm777):

@mathmale you there?

OpenStudy (anonymous):

\[3x+12=\pm 18,x=?,?\]

OpenStudy (snowystorm777):

4,6 @surjithayer

OpenStudy (mathmale):

this is an absolute value problem, so it needs to be treated a bit differently. You now have |x+4|=6. How many solutions does that have? What does the presence of the absolute value operator, | | , tell you to do? Given |x+4|=6, obtain two separate equations as follows: +(x+4)=6 -(x+4)=6 Solve these separately. check each result in the original equation.

OpenStudy (mathmale):

You could also try checking your supposed results (4 and 6) in the original equation.

OpenStudy (snowystorm777):

ok I am confused here @surjithayer can you finish your explantion

OpenStudy (anonymous):

3x+12=18,find x 3x+12=-18,find x

OpenStudy (mathmale):

Again, an absolute value problem such as yours can and should be broken up into two parts. surjithayer's approach and mine differ only in that I 've suggested dividing the entire equation through by 3 to simplify it, and I put my + and - signs on the left, whereas he put them on the right. the solutions are identical. Unfortunately, the solutions are not {4, 6}. If this is still not clear to you, then explain what you need to know.

OpenStudy (snowystorm777):

x+4=6 x+4=-6 ?? @surjithayer

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

find x for both equations.

OpenStudy (snowystorm777):

x=2 x=-2 @surjithayer

OpenStudy (anonymous):

no,x=-4-6=

OpenStudy (mathmale):

Please show your work. One of those is correct; the other is not.

OpenStudy (mathmale):

If you show your work, you'll get the best quality feedback.

OpenStudy (snowystorm777):

so x=-10 @surjithayer

OpenStudy (anonymous):

correct x=2,-10

OpenStudy (snowystorm777):

here is the choices A.No solution B.One solution C.Two solutions D.Infinitely many solutions SO THE ANSWER IS B TWO SOLUTIONS?

OpenStudy (snowystorm777):

@surjithayer

OpenStudy (anonymous):

how many values of x you got?

OpenStudy (anonymous):

it is c ,two solutions.

OpenStudy (snowystorm777):

OK THANKS!